L1 has y-intercept (0,4) and gradient -1/2
A. Find the equation of L1 in general form

Answer by mananth(16946)

y intercept 4
slope =-1/2
The equation in slope form
y = (-1/2)x +4
2y=-x +8
x+2y=8 is the equation